OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#8719 — DSLAM OVH Paris
Scheduled Maintenance Report for Web Cloud
Completed
An update is going live on the Paris DSLAMs. Works will begin at 00:00. Downtime of 30 minutes is to be expected on all sites from 00:30. We will be updating this task regularly in order to keep you well informed.

This new version brings some new functions:

- change of the DHCP lease from 5 mins to 1 week. The new MACs will be taken into account 2 mins rather than 5 mins before.
This means that in the event of changing modem or the router behind the modem in bridge mode, the DHCP server will transmit
the configuration faster.
- For multi-pair SDSLs with asymmetry, a function enabling pairs that disrupt others to be stablised and to thus have
the same flow on all pairs.

These functions will be activated later. The first stage is updating the DSLAMs.

Update(s):

Date: 2013-06-05 23:48:18 UTC
fla93-1-dsl, rny93-1-dsl, bny93-1-dsl, pla93-1-dsl, cou93-1-dsl, sto93-1-dsl, crn93-1-dsl, std93-1-dsl, bdy93-1-dsl, vil93-1-dsl are up.

Date: 2013-06-05 23:48:05 UTC
doing fla93-1-dsl, rny93-1-dsl, bny93-1-dsl, pla93-1-dsl, cou93-1-dsl, sto93-1-dsl, crn93-1-dsl, std93-1-dsl, bdy93-1-dsl, vil93-1-dsl.

Date: 2013-06-05 23:47:57 UTC
men75-1-dsl, bgl93-1-dsl, gam93-1-dsl, mtl93-1-dsl, dau94-1-dsl, fny94-1-dsl are ok

Date: 2013-06-05 23:47:47 UTC
doing men75-1-dsl, bgl93-1-dsl, gam93-1-dsl, mtl93-1-dsl, dau94-1-dsl, fny94-1-dsl.

Date: 2013-06-05 23:47:39 UTC
doing alf94-1-dsl, bos94-1-dsl, cre94-1-dsl, suc94-1-dsl, cnv94-1-dsl, pom94-1-dsl, cpy94-1-dsl, gra94-1-dsl, dag94-1-dsl are ok

Date: 2013-06-05 23:47:27 UTC
doing alf94-1-dsl, bos94-1-dsl, cre94-1-dsl, suc94-1-dsl, cnv94-1-dsl, pom94-1-dsl, cpy94-1-dsl, gra94-1-dsl, dag94-1-dsl.

Date: 2013-06-05 23:47:18 UTC
bny92-1-dsl che78-1-dsl vau92-1-dsl vit94-1-dsl are up.

Date: 2013-06-05 23:47:07 UTC
We will start upgrading the Dslams bny92-1-dsl che78-1-dsl vau92-1-dsl vit94-1-dsl



Date: 2013-06-05 15:15:49 UTC
We are going to upgrade the rest of the DSLAMs tonight at around 0:00.

Date: 2013-05-30 23:29:13 UTC
We will postpone the MAJ of the 3rd series to Monday.


Date: 2013-05-30 01:36:07 UTC
obs92 fixed, we'll upgrade the rest of dslams tomorrow night.

Date: 2013-05-30 01:12:06 UTC
bob75 is ok, 1 board left on obs92.

Date: 2013-05-30 01:11:59 UTC
1 board of tru75 is ok, 2 boards left bob75 & obs92, tech on his way there.

Date: 2013-05-30 01:11:52 UTC
we got 3 boards failed to boot, tech on his way to site.
we are suspending the rest of the upgrade till tomorrow.

Date: 2013-05-29 23:01:57 UTC
first set ok, starting with the 2nd one...

Date: 2013-05-29 23:01:49 UTC
rebooting first set..

Date: 2013-05-29 23:01:40 UTC
saving conf...
Posted May 29, 2013 - 14:29 UTC